OpenStudy (adrynicoleb):

A toy company manufactures arcade games. They are marketing a new pinball machine to children. It is similar in size to the adult version of the same game. Both the adult and child models are shown below. If the perimeter of the adult pinball machine is 172 inches, what is the length, in inches of line segment G'A'?

OpenStudy (adrynicoleb):

Here's the models.

OpenStudy (adrynicoleb):

@jdoe0001

OpenStudy (anonymous):

Well, we need to set up a proportion, correct? Use two similar sides, and set up a side first.

OpenStudy (adrynicoleb):

35/14?

OpenStudy (anonymous):

Alright, now we have to set up the second proportion. What do you think it will be?

OpenStudy (adrynicoleb):

56/x?

OpenStudy (anonymous):

You could, if you wanted. I believe you could also use the perimeter to find G' to A'

OpenStudy (adrynicoleb):

or 172/x?

OpenStudy (anonymous):

Honestly, I would find it on the big one first, it'll be less work.

OpenStudy (adrynicoleb):

ok so hold on a sec\

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Oh, i'm sorry! I meant find the length of G to A

OpenStudy (adrynicoleb):

this is probably wrong because it seems kinda high but i got 430

OpenStudy (anonymous):

definitely wrong, because we are finding the value of a smaller figure.

OpenStudy (anonymous):

So, let's find G to A. That's our total perimeter minus our known values, because we are only missing one side.

OpenStudy (adrynicoleb):

ok so i would do 172 - 35?? im confused

OpenStudy (anonymous):

Sorry that I cut out there, anyways, perimeter is the sum of all the sides, correct? Well, say x = length of segment AG. 35 + 56 + 56 + x = 172. Do you understand now?

OpenStudy (adrynicoleb):

yeah i think so. so i do 35 + 56 + 56, and then subtract it from 172??

OpenStudy (anonymous):

Exactly!

OpenStudy (adrynicoleb):

ok and then i get 25 right?

OpenStudy (anonymous):

Yes, now we can set up the second side of our proportion.

OpenStudy (anonymous):

Reminder, our first was \[\frac{ 35 }{14 }\]

OpenStudy (adrynicoleb):

ok so then it would be x/25? or the other way around?

OpenStudy (anonymous):

Remember, we but the adult pinball table on top, so the other way around.

OpenStudy (anonymous):

\[\frac{ 35 }{14} = \frac{ 25 }{ x }\]

OpenStudy (adrynicoleb):

ohhh i remember now. its 25/x because it is that way on the first side. got it. im gonna try to solve it hold on.

OpenStudy (adrynicoleb):

10??

OpenStudy (anonymous):

Yep! Solved.

OpenStudy (adrynicoleb):

YAY!! thank you SOOO much!! :D :D